OpenStudy (anonymous):

With the distance formula, I have seen that instead of people using d = square root of [ (X2 - X1)^2 + (Y2 - Y1)^2], they do this: d^2 = (X2 - X1)^2 + (Y2 - Y1)^2 ^ So they square both sides, and then they say that let S = d^2, meaning let this be the distance. Why is this the same as the first?

OpenStudy (anonymous):

it is not the distance, it is the square of the distance, but it is usually easier to work with with because there are no annoying square roots

OpenStudy (anonymous):

a calculus problem for example that would ask "what is the point closest to...?" it would be easier to work with the square of the distance than the distance

OpenStudy (anonymous):

d = square root of [ (X2 - X1)^2 + (Y2 - Y1)^2], they do this: d^2 = (X2 - X1)^2 + (Y2 - Y1)^2 squaring both side then u got confused because S=d^2 right? then u can write S=(X2 - X1)^2 + (Y2 - Y1)^2 it is just they supposed s as d^2